The number of photons emitted in 10 hours by a 60 W sodium lamp
(wavelength of photon = 6000 A)
    (A)    6.50 × 1024
    (B)    6.40 × 1023
    (C)    8.40 × 1023
    (D)    3.40 × 1023
    [Take hc = 12400 eVA, h = Planck’s constant, c = speed of light]

Grade:upto college level

1 Answers

Aditi Chauhan
10 years ago

(A)

 

Sol.   Let n be the number of photons emitted

 

 =>      (12400× 1.6 × 10–19 × n)/6000 = 60 × 10 × 60 × 6

 

        =>     n = 6.5 × 1024